OpenStudy (anonymous):

a(t) = A + Bt x(0) =xi v(0) = vi derive the position as a function of time equation

OpenStudy (zzr0ck3r):

take the first integral v(t) = At+(1/2)Bt^2 + C then plug in v(0) to solve for C

OpenStudy (3psilon):

Then take the next integral and plug X(0)

OpenStudy (zzr0ck3r):

yerp

OpenStudy (anonymous):

i take the integral of the first integral? and then plug in x(0)?

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

note you need to solve for c first

OpenStudy (anonymous):

how did you get (1/2)bt^2 ?

OpenStudy (3psilon):

anti derivative

OpenStudy (zzr0ck3r):

the integral of Bt

OpenStudy (zzr0ck3r):

A*0+(1/2)B*0^2 + C = vo c = v0

OpenStudy (anonymous):

ohhhhhh. okay okay. i get it now @zzr0ck3r

OpenStudy (zzr0ck3r):

great... note: Newton was a p i m p

OpenStudy (3psilon):

\[V(0) = V_{i} \] according to the boundary condition So \[V(0) =\frac{ B0^{2} }{ 2 } + a(0) + C\] \[ V(0) = C\] \[ V(0) = vi\] So \[ Vi = C\] Plug \[ V(i)\] into the orginal \[V(t)\] function as C